OpenStudy (fuzzyball):

Which statement is true? A. |–4| < 4 B. |–1| < 0 C. |–6| = –6 D. |–14| > 10

OpenStudy (fuzzyball):

I will fan and medal

OpenStudy (misty1212):

HI!!

OpenStudy (mathstudent55):

What is |-4| ?

OpenStudy (misty1212):

you want to make a guess?

OpenStudy (fuzzyball):

A

OpenStudy (mathstudent55):

Do you know what absolute value is?

OpenStudy (misty1212):

ok answer @mathstudent55 question first

OpenStudy (mathstudent55):

|0| = 0 |positive number| = positive number |negative number| = positive number

OpenStudy (fuzzyball):

no i dont

OpenStudy (mathstudent55):

The absolute value of a number is just the number with no sign. If the number was positive or zero, the absolute value is just the number. If the number was negative, the absolute value is just the number without the negative sign, so it is a positive number.

OpenStudy (anonymous):

The lines represent the absolute value. So for instance, if you had thr number -5, the absolute value of -5 would be 5 since absolute values cannot be negative.

OpenStudy (mathstudent55):

Examples: |1| = 1 |0| = 0 |-3| = 3 |-25| = 25 \(|\sqrt 2| = \sqrt 2\) \(|-\sqrt 3 | = \sqrt 3\)

OpenStudy (anonymous):

@mathstudent55 is correct.

OpenStudy (mathstudent55):

Do you understand the absolute value now?

OpenStudy (fuzzyball):

I think so..

OpenStudy (mathstudent55):

Good. Now look at choice A. |-4| < 4 We are comparing |-4| with 4 What is |-4| equal to?

OpenStudy (fuzzyball):

4?

OpenStudy (mathstudent55):

Correct. That means the statement in A is really 4 < 4 Is 4 less than 4?

OpenStudy (fuzzyball):

no

OpenStudy (mathstudent55):

Correct. That means statement A is false. We move on to B.

OpenStudy (mathstudent55):

B. |–1| < 0 What is |-1| ?

OpenStudy (mathstudent55):

Correct. A. 1 < 1 Is that true or false?

OpenStudy (fuzzyball):

false

OpenStudy (mathstudent55):

Correct. Move on to C.

OpenStudy (fuzzyball):

C is true because 6=6

OpenStudy (mathstudent55):

C. |–6| = –6 What is |-6| = ?

OpenStudy (mathstudent55):

Not so fast. What is |-6| = ?

OpenStudy (fuzzyball):

6

OpenStudy (mathstudent55):

That means C really is: |–6| = –6 6 = -6 Is 6 equal to -6?

OpenStudy (fuzzyball):

OH no

OpenStudy (mathstudent55):

Correct. By the process of elimination, the answer has to be D, but since we are not lazy, and we are here to learn, we will look at D.

OpenStudy (mathstudent55):

D. |–14| > 10 What is |-14| = ?

OpenStudy (fuzzyball):

14

OpenStudy (mathstudent55):

Correct. D is: |–14| > 10 14 > 10 Is 14 greater than 10?

OpenStudy (fuzzyball):

yes

OpenStudy (fuzzyball):

Thank you so much for your help

OpenStudy (mathstudent55):

Excellent. D is the correct answer since it is the only true statement among the choices.

OpenStudy (mathstudent55):

Great job!

OpenStudy (fuzzyball):

can you help me with a few more and check one i have already done?

OpenStudy (mathstudent55):

Yes, I have about 10 minutes I can help

OpenStudy (fuzzyball):

ok

OpenStudy (fuzzyball):

What is |–12|? i put 12

OpenStudy (mathstudent55):

Correct

OpenStudy (fuzzyball):

Which list of values is ordered from least to greatest? A. |–5|, 0, 3, –|–4| B. –7, –|–3|, 0, |–4| C. –|6|, |–3|, 1, 4 D. |–12|, –|–5|, 2, |–8|

OpenStudy (mathstudent55):

Remember, the absolute value is only the number without a sign, so if the number is zero or positive, it's just the number. If the number is negative, just take out the negative sign and make it positive.

OpenStudy (mathstudent55):

Take each choice and rewrite it by writing the correct absolute value. Here is A. A. |–5|, 0, 3, –|–4| 5, 0, 3, -4 |-5| = 5. You understand this, I'm sure. Do you understand why -|-4| is -4?

OpenStudy (fuzzyball):

yes because -I-4I it has and extra -

OpenStudy (mathstudent55):